### 17 Mar 2026 v2

# Index and Robustness of Mixed Equilibria: an algebraic approach

Abstract.
We present a new method for computation of the index of completely mixed equilibria in finite games, based on the work of Eisenbud et al. (1977). We apply this method to solving two questions about the relation of the index of equilibria and the index of fixed points, and the index of equilibria and payoff-robustness: any integer can be the index of an isolated completely mixed equilibrium of a finite game. In a particular class of isolated completely mixed equilibria, called monogenic, the index can be 0, +1+1 or 1-1 only. In this class non-zero index is equivalent to payoff-robustness. We also discuss extensions of the method of computation to extensive-form games, and cases where the equilibria might be located on the boundary of the strategy set.
I am grateful to Srihari Govindan and Rida Laraki for comments and stimulating conversations. I thank Julien Fixary for comments and proofreading.
1. Introduction
Topological methods are an integral part of game theory. Nash used the Brouwer fixed point theorem (Nash (1951)) and its generalization to correspondences (Nash (1950)) to prove existence of Nash equilibria, establishing a useful method that is widely used until this day. In refinements, most known solution concepts require robustness of equilibria to some kind of perturbation, and these have also been investigated by the use of topological tools. The stable equilibrium program, originated with work of Kohlberg and Mertens (1986), and the literature on dynamic stability in game theory (see Ritzberger (1994)) are exemplary in the modern theory of refinements in showcasing the centrality of topological methods for identifying equilibria that satisfy a number of desirable properties.
One of the cleanest criteria for equilibrium selection stemming out of this literature is the non-zero index rule. [^1] The index is an integer number associated to an equilibrium that captures the robustness of the equilibrium to perturbations of its associated best-reply correspondence to nearby continuous maps. Whenever the number is non-zero, the equilibrium is robust in this sense. Non-zero index equilibria are in particular robust to payoff perturbations, and are, for example, proper. Non-zero index equilibria have recently been axiomatized, for generic-games in extensive form, by Govindan and Wilson (2025) [^2] The axiomatization is of decision-theoretic nature and allows for a direct comparison to the solution concept of Mertens stability, also axiomatized in the same paper., and Pahl and Pimienta (2025) have proved a simple characterization that shows that these equilibria are hyperstable (Kohlberg and Mertens (1986)), i.e., they are the invariant and payoff-robust ones. [^3] Invariance here means that if we add finitely many duplicate strategies to the game, that is, mixed strategies as pure strategies, the hyperstable equilibria are still payoff-robust in the game with added strategies. Payoff-robustness of an equilibrium means that for sufficiently small payoff-perturbations, there are equilibria close to the given equilibrium point. In particular, in terms of desirable properties, non-zero index equilibrium outcomes (i.e., isolated equilibrium outcomes induced by a non-zero index component of mixed strategies) satisfy all the properties highlighted in Kohlberg and Mertens (1986): existence, admissibility, invariance, iterated dominance and sequential rationality. Because of these desirable properties, which are not simultaneously satisfied by other classical refinement concepts (perfect, proper, sequential equilibria, etc.), there is a strong case for the selection and computation of non-zero index equilibrium outcomes in finite games.
Unfortunately, in most non-trivial cases, the computation of the index is not an easy task to undertake. The topological tools involved in the computation of the index of equilibria do not fully explore the fact that the equilibria of finite games are described by polynomial systems of equalities and inequalities. In fact, when considering the standard formulae for computing the index of equilibria (see subsection 3.6 of Pahl (2023)), the particular polynomial nature of the maps is ignored. There are many difficulties that arise from this. An example: a known method to prove that a certain equilibrium is payoff robust is the verification of its non-zero index (this is a sufficient condition for payoff-robustness): to verify that, one can perturb the payoffs (in normal form) of the game to a nearby generic payoff function. Only a finite number of equilibria (with index +1 or -1) arise around the original equilibrium, and the index of this equilibrium is the sum of the indices of the equilibria that arose from the perturbation. There are a number of difficulties with the application of this procedure, for instance: how small should the generic payoff-perturbation be? Is there a canonical way to perturb the game in examples? After perturbing the game, is there a generally applicable method to find all equilibria nearby?
These questions exemplify one of the central issues with the application of the index of equilibria: the methods involve perturbations, and since there is no canonical way to handle these in general, the success of the method depends on clever choices made by the analyst in the specific example under analysis.
This paper presents a perturbation-free method to compute the index of isolated and completely mixed equilibria in finite games and applies it to solve a few questions about the relation of the index of fixed points and the index of equilibria: the first question is about the possible indices that isolated equilibria could have. In fixed point theory, isolated fixed-points can have any integer as an index. In game theory, it is known that if non-degenerate connected components of equilibria are allowed [^4] It is frequent that extensive-form games with generic terminal payoffs have equilibrium outcomes induced by a connected set of mixed strategies of the players - these are frequently non-singleton connected components of equilibria. In index theory, we assign an index to a component of equilibria., then these could have any integers as indices as well (Govindan et al. (2004)). However, the polynomial system describing isolated completely mixed equilibria seems quite particular - they are made of multilinear polynomials, and have the same number of variables as equations, thus suggesting, at least naively, a picture that would not allow for a wide variety of indices (in bimatrix games, the fact that an equilibrium is completely mixed forces its index to be 1-1 or +1+1, due to the linearity of the payoffs in the opponent’s strategy). Another reason the completely-mixed equilibrium environment under study looks restrictive comes from the main result in Govindan et al. (2025): the authors in that paper showed that - modulo the addition of duplicate strategies - any finite set of completely mixed strategy profiles, each with a corresponding sign of either +1+1 or 1-1, around a connected component of equilibria could be generated as equilibria of an arbitrarily close-by game (i.e., by an arbitrarily small payoff perturbation), as long as the sum of those signs is equal to the index of the component. This result seemed to indicate that completely mixed and isolated equilibria could only have indices of +1+1 or 1-1.
In this paper we show this is not the case: isolated completely mixed equilibria can be of any index. But the result is more nuanced: in a class of game-equilibrium pairs we call monogenic, and in which the isolated completely mixed equilibria are not too singular (i.e., where the Jacobian of the system describing it loses full rank by at most $1$ at this equilibrium point), the only possible indices for the equilibria under study are $0,+1$ and 1-1. More interestingly, in this class, the non-zero index criterion is equivalent to payoff-robustness: so, if in a more general context one needs to consider the invariance to addition of duplicates together with payoff-robustness to characterize non-zero index equilibria (see Govindan and Wilson (2005) and Pahl and Pimienta (2025)), in the monogenic environment payoff-robustness means non-zero index. Finally, we show in Theorem 3.13 that outside of the monogenic environment, isolated completely mixed equilibria can have any index.
These results have immediate implications for the refinement literature: it is known (see Example 3.11) that completely mixed equilibrium might not be robust to small payoff perturbations. The practical relevance of this observation is that the lack of robustness can only be ignored by the analyst if he/she believes that the modeled payoffs match exactly the payoffs in reality, thus making the concern of small parameter misspecification irrelevant. Since it is frequent that this is not the case, verifying payoff-robustness becomes an important selection criterion for equilibrium outcomes, and distinguishing between robust and non-robust equilibria is necessary. Directly verifying payoff robustness is usually impractical even in simple examples, thus making the index a valuable tool. But the non-zero index criterion in general is known to merely imply payoff-robustness. As we show, in the monogenic case, the non-zero index criterion is also necessary for payoff robustness. Moreover, the algebraic method employed to verify the non-zero index of equilibria can be significantly easier to apply than the standard available methods.
This paper is organized as follows: section 2 comprises preliminary facts on algebra and topology that we use throughout this paper: we present a number of definitions in basic abstract algebra and algebraic geometry that might not be in the toolkit of the theoretical economist and game theorist and also refer to the relevant literature for consultation. We also prove the main index formula that will be used throughout the paper (Proposition 2.1). Section 3 presents the main results on the monogenic class and its relevance for payoff-robustness (Theorem 3.4). We also present a detailed application of the index formula for computation of the index of an equilibrium in an example (Example 3.11), and the result stating that isolated completely mixed equilibria can have any index (Theorem 3.13). In Section 4 we discuss extensions of the main results to extensive-form games, to the case of non-degenerate components of equilibria, and equilibria that might not be completely mixed. Section 5 contains final remarks on open problems and a discussion of the cases to which the results in this paper do not apply.
1.1. A motivating example
We show through the analysis of an example, due to E. Solan and O.N. Solan in McLennan (2018), p. 151, how verifying the sign of the index of an equilibrium can be a complicated task with the use of standard tools. Later, we revisit this example to show how index computation is straightforward applying the new tools. Consider the following three-player game:
$a$ $b$
$a$ $(1,1,1)$ (5,0,3)(-5,0,3)
$b$ (0,3,5)(0,3,-5) $(0,0,1)$
Player 3 chooses $a$
$a$ $b$
$a$ (3,5,0)(3,-5,0) $(1,0,0)$
$b$ $(0,1,0)$ $(0,0,0)$
Player 3 chooses $b$
There is a completely mixed equilibrium $\sigma^{*}$ in this game, where all players use equal weights for all strategies. The simplest way to compute the index of this equilibrium is to attempt to use some of the index properties. In this case that the sum of the indices of all equilbrium connected components must be +1+1. But this immediately imposes the difficult task of checking whether the game has other equilibrium points. If it does, then, hopefully, other properties of the index of equilibria could eventually be used to pin down the sign of the index of $\sigma^{*}$. In the example, it is easy to see that there is a strict equilibrium in $(a,a,a)$, and index theory tells us that such an equilibrium must have index +1+1. Therefore, all other equilibria of the game must have indices whose sum is 0. If somehow one could check that other equilibria - besides the completely mixed - do not exist, we would conclude that $\sigma^{*}$ has index 0. But as mentioned, even in this game, this is not straightforward.
The next method one could use is to check whether the Jacobian of a system describing the completely mixed equilibrium of the game $\sigma^{*}$ is non-singular. The following system, obtained by payoff-indifference conditions, where we translate $\sigma^{*}$ to the point 0 would do:
(1) $p_{1}$ $=x-y+xy=0$
$p_{2}$ $=y-z+yz=0$
$p_{3}$ $=z-x+zx=0$
Verifying that the Jacobian of this system at 0 is non-singular would imply that the equilibrium has non-zero index. But a quick verification shows that the Jacobian at $\sigma^{*}$ is singular. These are the easiest methods to compute the sign of the index of $\sigma^{*}$. Another frequently applied method involves payoff perturbations: one could use the property that the index is locally constant with respect to payoff perturbations, that is, fixing a neighborhood of $\sigma^{*}$ in the mixed strategy set that contains no other equilibrium in its closure, there exists a $\delta>0$ such that for any generic $\delta$-perturbation of the payoffs of the game, there are finitely many equilibria in the interior of the neighborhood, each with an index of +1+1 or 1-1 summing to the index of $\sigma^{*}$. The difficulty with using this property is clear: one in principle doesn’t know how small $\delta>0$ must be, nor have any idea how many equilibria may arise in the neighborhood; one needs to determine all these equilibria and compute the index of each one separately.
If for some reason one believes $\sigma^{*}$ has index 0, one could attempt to construct a perturbation of the payoffs that has no nearby equilibrium (this is also not straightforward), which is, in general, only a sufficient condition for 0-index. We return to this example later to show how the index can be computed from analysing simple algebraic properties of the system of polynomials in (1).
2. Preliminaries
We review some basic algebraic and topological facts that are required for this paper. For the algebra, a general practical source for what is not defined in this paper is Cox et al. (2005). For the topological degree and index theories a standard reference is Dold (1972).
2.1. Algebraic Preliminaries
A monomial in a collection of $n$-variables $X_{1},....,X_{n}$ is a product $X^{\alpha}=X^{\alpha_{1}}_{1}...X^{\alpha_{n}}_{n}$, where $\alpha_{i}$ is a non-negative integer. Real polynomials in $n$ variables can be written as finite linear combinations of monomials, with real coefficients. Formal real power series comprise the set of infinite linear combinations of monomials with real coefficients. Denote by $\mathbb{R}[[X_{1},....,X_{n}]]$ the set of formal power series. The set of real formal power series is an example of a commutative ring with unity, where operations of addition and multiplication are the usual operations one generalizes from polynomial rings (which are also commutative rings with unity).
An ideal $I$ of a ring $R$ is a subset of $R$ that is closed for addition ($x,y\in I\implies x+y\in I$) and absorbs multiplication ($x\in R,y\in I\implies xy\in I$). An ideal $I$ is finitely generated if there exists $f_{1},...,f_{n}\in I$ such that for any element $x\in I$, there exist $a_{i}\in R,i=1,....,n$, such that $x=\sum_{i}a_{i}f_{i}$. In this case, we denote $I=\langle f_{1},...,f_{n}\rangle$. If $I,J$ are ideals of $R$, $I+J$ (the sums of elements of each ideal) and $I\cap J$ are ideals of $R$. For a finite positive integer $k$, $I^{k}$ denotes the smallest ideal generated by the product of $k$ elements of $I$. An ideal is called square-zero if $I^{2}=0$. An ideal that is generated by a single element is called principal.
Given an ideal $I\subset R$, say that we define an equivalence relation $\sim$ in $R$, by letting $x\sim y\iff x-y\in I$. The equivalence class of $x$ is denoted $[x]$. The set of equivalence classes under this relation is denoted $R/I$ or $\frac{R}{I}$, and is given a ring structure from the operations of $R$: $x,y\in R,[x]\oplus[y]:=[x+y]$ and $z,x\in R,[z]\odot[x]:=[zx]$. This ring defined on the equivalence classes is called a quotient ring. If $I\subset J$ are ideals of $R$, then $J/I$ is also defined as the set of equivalence classes of elements of $J$ and is an ideal of $R/I$.
In this paper we will mostly consider ideals of real formal power series that are finitely generated, and quotient rings of such ideals. Note that the quotient rings of real formal power series can also be given a real vector space structure from the ring operations on it.
[…]
Real polynomials have a natural real vector space structure: they are trivially closed by addition (we have already observed that from their ring structure), but also by scalar multiplication by an element of $\mathbb{R}$, thus satisfying the properties of an $\mathbb{R}$-algebra. By the same reason, the real formal power series are an $\mathbb{R}$-algebra as well. Therefore, they are both rings and real vector spaces. We say that two $\mathbb{R}$-algebras $R$ and $L$ are homomorphic if there exists an $\mathbb{R}$-linear homomorphism $\psi:R\to L$ such that $\psi(xy)=\psi(x)\psi(y)$. The homomorphism is unital if it takes the multiplication unity element of $R$ to the multiplication unity of $L$. All algebra-homomorphisms that appear in this paper are unital. An isomorphism of $\mathbb{R}$-algebras (i.e., an $\mathbb{R}$-algebra homomorphism with an inverse $\mathbb{R}$-algebra homomorphism) is also denoted by $\cong$. We will distinguish from just a ring-isomorphism by explicit mention to it. Note that an ideal $I$ of an $\mathbb{R}$-algebra that is closed by scalar multiplication is also a linear subspace of this algebra.
The ring of real formal power series is an example of a local ring: a ring that has exactly one maximal ideal (a maximal ideal is a proper ideal of a ring that is not properly contained in any other ideal). The maximal ideal of $\mathbb{R}[[X_{1},....,X_{n}]]$ is the finitely generated ideal $\mathfrak{m}=\langle X_{1},...,X_{n}\rangle$. Note that if $f\notin\mathfrak{m}$, then $f$ can be inverted - i.e., one can explicitly write down a real formal power series $g$ which is the multiplication inverse of $f$. This simple fact is crucial for solving Example 3.11.
If a list of $m$ real formal power series in $n$ variables $p_{1},....,p_{m}$ is such that for each $i$, $p_{i}$ has no constant term, then it follows that $\langle p_{1},...,p_{m}\rangle\subset\mathfrak{m}$. In particular, if for each $i$, $p_{i}=0$ is a system of polynomials with 0 as an isolated root, then each $p_{i}$ has no constant term. Therefore, $\langle p_{1},....,p_{m}\rangle\subset\mathfrak{m}$.
For an ideal $I\subset\mathfrak{m}\subset\mathbb{R}[[X_{1},....,X_{n}]]$, $\mathbb{R}[[X_{1},...,X_{n}]]/I$ is a ring, but also inherits a real vector space structure from $\mathbb{R}[[X_{1},...,X_{n}]]$, which makes it into an $\mathbb{R}$-algebra as well. Determining which class $[f]$ is non-zero in this quotient is equivalent to determining if $f\in I$. This is relevant, for example, if one wants to know if a linear combination in $\mathbb{R}[[X_{1},...,X_{n}]]/I$ is zero, or if one wants to construct a basis for the vector space $\mathbb{R}[[X_{1},...,X_{n}]]/I$ in order to establish its dimension, something that is intimately linked to the problem of computing the index of an isolated equilibrium (see Proposition 2.1). For our goal in this paper, we will be interested in the case where $I$ is finitely generated by polynomials $p_{1},....,p_{m}$. Therefore we need to establish that there exists $a_{i}\in\mathbb{R}[[X_{1},...,X_{n}]],i=1,...,n$ such that $f=\sum_{i}a_{i}p_{i}$. Mora’s normal form algorithm allows us to answer this question, by generalizing the known division algorithms for polynomials (see Cox et al. (2005), ch.4, Corollary 3.14 and exercise 2, which shows one can perform division with formal power series as inputs). In order to do that, we need to define a local order on the set of monomials: an order >> on the set of monomials is a local order if it is total, compatible with multiplication ($\alpha,\beta,\gamma\in\mathbb{Z}^{n}_{+},X^{\alpha}>X^{\beta}\implies X^{\alpha+\gamma}>X^{\beta+\gamma}$) and for each $i=1,...,n,1>X_{i}$. A particular local order we use is called the degree-anticompatible lex order: in this order $\alpha,\beta\in\mathbb{Z}^{n}_{+},X^{\alpha}>_{a\ell ex}X^{\beta}$ if
$|\alpha|=\sum_{i}\alpha_{i}<|\beta|=\sum_{i}\beta_{i}$
or if
$|\alpha|=|\beta|\text{ and }X^{\alpha}>_{\ell ex}X^{\beta},$
where $_{\ell ex}$ is some lexicographic order on monomials.
With a local order, one can define the leading term of a formal power series: if $f=\sum_{\alpha\in\mathbb{Z}^{n}_{+}}c_{\alpha}X^{\alpha}$, then its leading term LT(f)(f) is the $c_{\bar{\alpha}}X^{\bar{\alpha}}$ where $X^{\bar{\alpha}}$ is maximal under the local order >> over the monomials defining $f$. Given an ideal $I\subset\mathbb{R}[[X_{1},...,X_{n}]]$, the leading term ideal of I $\langle LT[I]\rangle$ is the ideal generated by the leading terms of the power series in $I$. This is an example of a monomial ideal, an ideal that has a set of generators composed of monomials. The leading term ideal has a few crucial properties that allow us to quickly perform the verification of whether an index is 0 or not: the first is given by Theorem 4.3 in Cox et al. (2005), namely, when the dimension of $\mathbb{R}[[X_{1},...,X_{n}]]/I$ is finite, the dimension of $\mathbb{R}[[X_{1},...,X_{n}]]/I$ is equal to $\mathbb{R}[[X_{1},...,X_{n}]]/\langle LT[I]\rangle$. Verifying that a polynomial is a member of a monomial ideal is usually a simpler task than doing it when the ideal does not have such a property, and we will explore precisely this in Example 3.11. In addition, the dimension of $\mathbb{R}[[X_{1},...,X_{n}]]/\langle LT[I]\rangle$ is simply the number of standard monomials, i.e., those monomials $X^{\alpha}\notin\langle LT[I]\rangle$ (that is, monomials that are not divisible by any of the generators in $LT[I]$).
Note that when we consider real formal power series rings in one variable, there is only one local order ($1>X>X^{2}>.$...). The leading term of such a power series $f$ is then the term that has the smallest degree, also known as the order of $f$, denoted ord($f$).
2.2. Topological Preliminaries
The main tools that will be used in this paper are the index and degree of equilibria. These are both specifications to the environment of Nash equilibria of the concepts of topological degree and index (cf. Ritzberger (1994)). Section 2 in Pahl (2023) reviews precisely the information required for this paper, explaining intuitively the roles of the structure theorem of Kohlberg and Mertens (Theorem 1 in Kohlberg and Mertens (1986)), degree theory and index theory in the problem of identifying payoff-robust equilibria, as well as formally presenting the definitions. We do not think that repeating this presentation here would be significantly more helpful relative to the cost of making the paper longer. We refer to that paper for these preliminaries. Other standard references are McLennan (2018), Dold (1972) and Brown (2014).
Notationwise, the index of a fixed point $x$ of a map $f$ will be denoted ind(x,f)(x,f). The index of an equilibrium $\sigma^{*}$ refers to the index of the isolated equilibrium $\sigma^{*}$ of a game $G$ with respect to any Nash-map that has $\sigma^{*}$ as fixed point when the game $G$ is fixed (see subsection 2.2.2 in Pahl (2023) for details on Nash-maps). For an open set $U\subset\mathbb{R}^{m}$ such that the closure $\bar{U}$ contains a unique solution $x$ of the equation $p(x)=0\in\mathbb{R}^{m}$, we denote by $\text{deg}_{0}(x,p)$ the topological degree of x w.r.t. to $p$. When the point $x$ is understood, we omit reference to it and write $\text{deg}_{0}(p)$. Recall that the degree and the index of an isolated equilibrium are identical (the result actually holds for connected components of equilibria). For two players, the result is proved in Govindan and Wilson (1997). The general result is proved in DeMichelis and Germano (2000).
The next proposition provides a formula for the computation of the absolute value of the topological degree of an isolated root of a polynomial system, which is directly based on the formula obtained by Eisenbud and Levine in Theorem 1.1 in Eisenbud et al. (1977). In their formula, the $\mathbb{R}$-algebras considered are germs at 0 of smooth real maps and the condition for the formula to hold is that the germ is finite. To make the application of the formula straightforward, it is worth recasting the formula in terms of $\mathbb{R}$-algebras of formal power series, and the condition in terms of finite dimensionality of the $\mathbb{R}$-algebras. The proposition is certainly known, but we include a proof here for completeness. In the proof of Theorem 3.13, we employ another formula for the computation of the degree presented in Theorem 1.2 in Eisenbud et al. (1977).
Proposition 2.1.
Suppose $f:(\mathbb{R}^{n},0)\to(\mathbb{R}^{m},0)$ is a polynomial map with an isolated root at 0. Suppose $I=\langle f_{1},...,f_{m}\rangle$ is an ideal of $\mathbb{R}[[X_{1},....,X_{n}]]$ and $\mathbb{R}[[X_{1},....,X_{n}]]/I$ has finite dimension. Then
$|\text{deg}_{0}(0,f)|=\text{dim}_{\mathbb{R}}(\mathbb{R}[[X_{1},....,X_{n}]]/I)-2\text{dim}_{\mathbb{R}}I_{max},$
where $I_{max}$ is a maximal square-zero ideal of $\mathbb{R}[[X_{1},....,X_{n}]]/I$.
Proof.
Let $S=C^{\infty}_{0}(\mathbb{R}^{n})$ (the $\mathbb{R}$-algebra of germs of smooth functions at 0) and $A=\mathbb{R}[[X_{1},...,X_{n}]]$. We denote by $I_{S}$ the ideal the generated in $S$ by the $f_{i}$’s and analogously by $I_{A}$ the ideal generated in $A$ by the same $f_{i}$’s. Let $T:S\to A$ be the ring homomorphism associating to each smooth function germ $f$ at 0, its Taylor series at 0. This map is surjective, by Borel’s theorem. Hence $T$ induces a surjective map $\hat{T}:S\to\frac{A}{I_{A}}$ by composing $T$ with the partition map from $A$ to $\frac{A}{I_{A}}$. Now, since each generator $f_{i}$ is polynomial, $T(f_{i})=f_{i}$, which implies that $T(I_{S})\subset I_{A}$. Therefore, the induced map $\bar{T}:\frac{S}{I_{S}}\to\frac{A}{I_{A}}$ is still surjective. In order to prove that $\bar{T}$ is actually an isomorphism, we need to prove that Ker($\bar{T}$) $\{0\}$. To do that we establish the following lemma.
Lemma 2.2.
$Ker(\bar{T})$ is ring-isomorphic to $\frac{Ker(T)}{Ker(T)\cap I_{S}}$.
Proof of the Lemma.
[…]
Given Lemma 2.2, $Ker(\bar{T})\cong\frac{\text{Ker}(T)}{Ker(T)\cap I_{S}}=\{0\}$ if and only if $Ker(T)\subset I_{S}$.
We now establish that $Ker(T)\subset I_{S}$. First recall that Ker(T)(T) comprises the flat germs, that is, germs of smooth functions at 0 such that the derivatives at all orders vanish at 0. Also recall that $S$ is a local ring with maximal ideal denoted by $\mathfrak{m}_{S}$ generated by $X_{1},\ldots,X_{n}$. Using the known fact that
$Ker(T)\subset\bigcap_{k\geq 1}\mathfrak{m}^{k}_{S},$
it is sufficient to prove that there exists $k\in\mathbb{N}$ for which $\mathfrak{m}^{k}_{S}\subset I_{S}$. Let $R$ denote the localized ring of polynomials $\mathbb{R}[x_{1},...,x_{n}]_{\mathfrak{m}}$. By assumption, as $\text{dim}_{\mathbb{R}}\frac{A}{I_{A}}$ is finite, Buchberger’s algorithm and Theorem 4.3 in Cox et al. (2005), imply that $\frac{A}{I_{A}}$ and $\frac{R}{I_{R}}$ have the same dimension as real vector spaces and have the same standard monomials as well. This implies that there exists $k\in\mathbb{N}$ such that $\mathfrak{m}_{R}^{k}\subset I_{R}$. Let now $X^{\alpha}$ be a monomial such that $|\alpha|=k$. Recall that such monomials generate $\mathfrak{m}^{k}_{S}$ and $\mathfrak{m}^{k}_{R}$. Since $X^{\alpha}\in I_{R}$, then $X^{\alpha}=\frac{f}{g}$, $f\in J,g\in R\setminus\mathfrak{m}_{R}$, and $J\subset\mathbb{R}[X_{1},...,X_{n}]$ is the ideal generated by $f_{i},i=1,\ldots,m$. Therefore, $\frac{f}{g}$ can be written as $\sum^{m}_{i=1}\frac{a_{i}}{g}f_{i},a_{i}\in\mathbb{R}[x_{1},...,x_{n}]$. This implies $\frac{f}{g}\in I_{S}$, since $\frac{a_{i}}{g}$ represents a smooth function germ at 0. Hence, $\mathfrak{m}^{k}_{S}\subset I_{S}$ thus concluding the proof that $\bar{T}$ is a ring-isomorphism. Note that, in addition, $\bar{T}$ is $\mathbb{R}$-linear, so $\bar{T}$ is an isomorphism of $\mathbb{R}$-algebras. This implies that both $S/I_{S}$ and $A/I_{A}$ have the same dimension. In particular, the maximal square-zero ideal $J_{max}$ of $\frac{S}{I_{S}}$ maps under $\bar{T}$ to the maximal square-zero ideal $I_{max}$ of $\frac{A}{I_{A}}$, with both having the same dimension as well. Applying then Theorem 1.1 in Eisenbud et al. (1977), the result follows.∎
Remark 2.3.
If 0 is isolated as a complex root of $f$, then one can show that $\mathbb{R}[[X_{1},....,X_{n}]]/I$ has finite dimension - thus making the assumption of finite real-dimension of $\mathbb{R}[[X_{1},....,X_{n}]]/I$ redundant. However, if 0 is isolated as a real root, but not as a complex root, the assumption of finite dimension of $\mathbb{R}[[X_{1},....,X_{n}]]/I$ is still required to obtain Proposition 2.1: take as an example $f(x,y)=((x^{2}+y^{2})x,(x^{2}+y^{2})y)$. The point 0 is an isolated real root (but not as a complex root) and $\mathbb{R}[[X_{1},....,X_{n}]]/I$ has infinite dimension.
3. Index and Robustness
[…]
An isolated equilibrium $\sigma$ is called payoff-robust if for any neighborhood $U\subset\Sigma$ of $\sigma$, there exists a neighborhood $V\subset\mathbb{R}^{S}$ of $G$ such that for any $G^{\prime}\in V$ there exists an equilibrium $\sigma^{\prime}\in U$ of $G^{\prime}$. Non-zero index and non-zero degree equilibria are immediately robust to payoff perturbations. This is due to Ritzberger (1994) (see section 2 in Pahl (2023) for a discussion).
We need a specific result on the degree and index of equilibria for the proof of part (b) of Theorem 3.4. Let $r_{n}:\mathbb{R}^{S_{n}}\to\Sigma_{n}$ denote the closest-point retraction from a point $z_{n}\in\mathbb{R}^{S_{n}}$ to a point in $\Sigma_{n}$, and let $r=\times_{n}r_{n}$. Letting $\mathcal{E}_{G}=\{(g,\sigma)\in\mathbb{R}^{\kappa+N}\times\Sigma\mid\sigma\text{ is an equilibrium of }G\oplus g\}$, where $G\oplus g$ is a finite game, with the mixed strategy set of each player $n$ as $\Sigma_{n}$ and payoffs defined as $G_{n}(\sigma)+g_{n}$. Define $\theta_{G}:\mathcal{E}_{G}\to\mathbb{R}^{\kappa+N}$ by $\theta_{G}(g,\sigma)=(\sigma_{n}+G^{n}(\sigma)+g_{n})_{n\in\mathcal{N}}$. Using a similar argument as in Kohlberg and Mertens, one can obtain that $\theta_{G}$ is a homeomorphism, with inverse $\theta_{G}^{-1}(z)=(g(z),r(z))$, where $g_{n}(z)=z_{n}-r_{n}(z_{n})-G^{n}(r_{-n}(z_{-n}))$, and $g=\times_{n}g_{n}$. Letting proj: $\mathcal{E}\to\mathbb{R}^{\kappa+N}$ denote the projection over the payoffs $g$, for each isolated equilibrium $\sigma^{*}$ of $G$, and $U\subset\mathbb{R}^{\kappa+N}$ a neighborhood of $\theta_{G}(0,\sigma^{*})=z^{*}$ containing in its closure no other $z$ such that $\theta_{G}(0,\sigma)=z$, the topological degree over 0 of $\text{proj}\circ\theta^{-1}_{G}:U\to\mathbb{R}^{\kappa+N}$ is the index of $\sigma^{*}$ (see Govindan and Wilson (2005) Formulation and Appendix A for a detailed explanation of this fact).
Proposition 3.1.
Given any finite game $G$ with $\sigma^{*}$ a completely mixed and isolated equilibrium of $G$, it follows that the topological degree of 0 with respect to $p$ equals the index of $\sigma^{*}$.
Proof.
Define
$f_{s_{n}}(\sigma)=\frac{\sigma_{s_{n}}+G^{n}(\sigma)\cdot(e_{s_{n}}-e_{s^{0}_{n}})}{1+\sum_{\tilde{s}_{n}\in S_{n}}G^{n}(\sigma)\cdot(e_{\tilde{s}_{n}}-e_{s^{0}_{n}})}.$
[…]
Let
$h_{s_{n}}(\sigma)=\frac{\sigma_{s_{n}}+\text{max}\{0,G^{n}(\sigma)\cdot(e_{s_{n}}-\sigma_{n})\}}{1+\sum_{\tilde{s}_{n}\in S_{n}}\text{max}\{0,G^{n}(\sigma)\cdot(e_{\tilde{s}_{n}}-\sigma_{n})\}},$
with $h_{n}=\times h_{s_{n}}:\Sigma\to\Sigma$, and $h=\times_{n}h_{n}$, be the map used in Nash (1951) to prove existence of Nash equilibria. In what follows, a homotopy $H:[0,1]\times U\to\Sigma$ between $H(0,\cdot)=f:U\to\Sigma$ and $H(1,\cdot)=h:U\to\Sigma$ is called admissible if the boundary of $U$ contains no fixed points of the homotopy, i.e., for each $\sigma\in\partial U$ and $t\in[0,1]$, $H(t,\sigma)\neq\sigma$.
Lemma 3.2.
There is an admissible homotopy between $f:U\to\Sigma$ and $h:U\to\Sigma$. Therefore, $\text{ind}(\sigma^{*},f)=\text{ind}(\sigma^{*},h)$.
Proof of the Lemma.
Let $\phi_{s_{n}}(\sigma)=G^{n}(\sigma)\cdot(e_{s_{n}}-e_{s^{0}_{n}})$ and $q_{s_{n}}(\sigma)=\text{max}\{0,G^{n}(\sigma)\cdot(e_{s_{n}}-\sigma_{n})\}$. Adjusting the neighborhood $U$ if necessary to a smaller one around $\sigma^{*}$, define the homotopy $H:[0,1]\times U\to\Sigma$,
$H_{s_{n}}(t,\sigma)=\frac{\sigma_{s_{n}}+(1-t)\phi_{s_{n}}(\sigma)+tq_{s_{n}}(\sigma)}{1+\sum_{s_{n}\in S_{n}}(1-t)\phi_{s_{n}}(\sigma)+tq_{s_{n}}(\sigma)}.$
Let $0<t<1$. Suppose by contradiction $\sigma$ is a fixed point of the homotopy distinct from $\sigma^{*}$. Therefore, $\sigma$ is not a Nash-equilibrium of $G$.
Fix $n\in\mathcal{N}$ such that $\sigma_{n}$ is not a best-reply to $\sigma$. Suppose first $s^{0}_{n}$ is a best-reply to $\sigma$ in $G$. Then $\phi_{s^{0}_{n}}(\sigma)=0$ and $q_{s^{0}_{n}}(\sigma)>0$. Therefore, $\sum_{s_{n}}(1-t)\phi_{s_{n}}(\sigma)+tq_{s_{n}}(\sigma)>0.$ There exists now $\bar{s}_{n}\in S_{n}$ a pure strategy that yields an inferior payoff against $\sigma_{-n}$ than $\sigma_{n}$ (since $\sigma_{n}$ is completely mixed). Therefore, $\phi_{\bar{s}_{n}}(\sigma)<0$ and $q_{\bar{s}_{n}}(\sigma)=0$, thus implying $\sum_{s_{n}}(1-t)\phi_{s_{n}}(\sigma)+tq_{s_{n}}(\sigma)<0,$ which is a contradiction. Therefore, $s^{0}_{n}$ cannot be a best-reply to $\sigma$. Let now $\hat{s}_{n}$ be a best-reply to $\sigma$, which implies that $\sum_{s_{n}}(1-t)\phi_{s_{n}}(\sigma)+tq_{s_{n}}(\sigma)>0.$ Let now $\tilde{s}_{n}$ yield a payoff against $\sigma$ that is at most as good as $s^{0}_{n}$ and inferior to $\sigma_{n}$. This implies then $\sum_{s_{n}}(1-t)\phi_{s_{n}}(\sigma)+tq_{s_{n}}(\sigma)<0$. Contradiction. Therefore, the only fixed point of $H(t,\cdot)$ is $\sigma^{*}$ and the homotopy is in particular admissible.∎
Let $A^{0}_{n}=\{X_{n}=(X_{s_{n}})_{s_{n}\in S_{n}}\mid\text{ for each $n\in\mathcal{N}$, }\sum_{s_{n}\in S_{n}}X_{s_{n}}=0\}$, and $A^{0}=\times_{n}A^{0}_{n}$. Let $F_{n}$ denote the affine space generated by $\Sigma_{n}$. Consider the homeomorphism $T_{n}:A^{0}_{n}\to F_{n}$, where $T_{n}=\times_{s_{n}}T_{n}$, $T_{s_{n}}(X_{n})=X_{s_{n}}+\sigma^{*}_{s_{n}}$. Let $V=T^{-1}(U)$. Let $T=\times_{n}T_{n}$. Define $\hat{f}=T^{-1}\circ f\circ T:V\to A^{0}_{n}$. By the commutativity property of the index, $\text{ind}(0,\hat{f})=\text{ind}(\sigma^{*},f)$. Note now that for each $s_{n}\in S_{n}$,
$\hat{f}_{s_{n}}(X)=\frac{T_{s_{n}}(X)+G^{n}(T(X))\cdot(e_{s_{n}}-e_{s^{0}_{n}})}{1+\sum_{\tilde{s}_{n}\in S_{n}}[G^{n}(T(X))\cdot(e_{\tilde{s}_{n}}-e_{s^{0}_{n}})]}-\sigma^{*}_{s_{n}}.$
Consider now the map $S:\mathbb{R}^{\kappa}\to A^{0}$ given by $S=\times_{n}S_{n}$ with
$S_{n}(X_{s^{1}_{n}},\ldots,X_{s^{k_{n}}_{n}})=(-\sum^{k_{n}}_{i=1}X_{s^{i}_{n}},X_{s^{1}_{n}},\ldots,X_{s^{k_{n}}_{n}}).$
[…]
$Q_{s_{n}}(t,X)=\frac{T_{s_{n}}(X)+G^{n}(T(X))\cdot(e_{s_{n}}-e_{s^{0}_{n}})}{1+(1-t)\sum_{s_{n}\in S_{n}}[G^{n}(T(X))\cdot(e_{s_{n}}-e_{s^{0}_{n}})]}-\sigma^{*}_{s_{n}}.$
A similar argument to the one in the proof of Lemma 3.2 yields that 0 is the only fixed point of the homotopy which then implies that $(0,\hat{h})= ind$$(0,Q(1,\cdot)$). Now, notice that $p_{s_{n}}(X)=X_{s_{n}}-Q_{s_{n}}(1,\cdot)$, so the displacement of $Q(1,\cdot)$ equals $p$. Since the index of 0 w.r.t. $Q(1,\cdot)$ equals the topological degree of 0 w.r.t. its displacement (by definition), then the result follows. ∎
Definition 3.3.
The game-equilibrium pair $(G,\sigma^{*})$ is monogenic if $\sigma^{*}$ is an isolated completly mixed equilibrium of $G$ and $Jp(0)$ has rank $\geq\kappa-1$.
The definition of a monogenic game-equilibrium pair encompasses nongeneric cases of isolated completely mixed equilibria, where the Jacobian matrix of $p$ might lose full rank by at most one at the root. Generically, in normal-form games, all equilibria are isolated and have indices +1+1 or 1-1 only. Extensive-form games (even with generic terminal payoffs) might have non-isolated equilibria (we discuss the application of our results in these cases in section 4). Theorem 3.4 shows that when the Jacobian loses full rank by at most one, only one new possibility of index arises for isolated completely mixed equilibria relative to the generic case, namely, the index 0 (this is result $(a)$ in Theorem 3.4). We also show in item $(b)$ of the same theorem that in the monogenic case, non-zero index fully characterizes payoff-robustness, sharpening the characterization of non-zero index components in Pahl and Pimienta (2025) and Govindan and Wilson (2005) on which we commented in the introduction.
Theorem 3.4.
Suppose $(G,\sigma^{*})$ is monogenic. The following statements hold:
  1. (a)
    The index of $\sigma^{*}$ is either $0,+1$ or 1-1;
  2. (b)
    $\sigma^{*}$ has index 0 if and only if $\sigma^{*}$ is not payoff-robust.
Proof.
We start with the proof of item (a) of the theorem. For that we prove an auxiliary lemma. Let $I=\langle p_{1},...,p_{\kappa}\rangle$ be the ideal generated by the polynomial map $p$ associated to $\sigma^{*}$.
Lemma 3.5.
The pair $(G,\sigma^{*})$ is monogenic if and only if there exists a one-variable real power series $g(w)$ with $g(0)=0$ such that $\mathbb{R}[[X_{1},....,X_{\kappa}]]/I\cong\mathbb{R}[[w]]/\langle g(w)\rangle$ as $\mathbb{R}$-algebras.
Notice that, provided the lemma holds, statement (a) is almost immediate: given the $\mathbb{R}$-algebra isomorphism expressed in the lemma, it follows immediately that $\mathbb{R}[[X_{1},....,X_{\kappa}]]/I$ is finite dimensional. By Proposition 2.1 we then have that $\text{deg}_{0}(p)=\text{dim}_{\mathbb{R}}\mathbb{R}[[w]]/\langle g(w)\rangle-2\text{dim}_{\mathbb{R}}I^{w}_{max}=\text{deg}_{0}(g)$, where $I^{w}_{max}$ is a maximal square-zero ideal in $\mathbb{R}[[w]]/\langle g(w)\rangle$. Now, it is known (see for example O’Neill (1953)) that in the case of one-dimensional real maps the local degrees can only be 0, +1+1 or 1-1, which concludes the proof of item (a).
Proof of Lemma 3.5.
Suppose $(G,\sigma^{*})$ is monogenic. After a change of variables, without loss of generality we can assume that the submatrix of the Jacobian of $p$ at 0 given by $\frac{\partial(p_{1},...,p_{\kappa-1})}{\partial(x_{1},...,x_{\kappa-1})}(0)$ is invertible. Define $\Psi:\mathbb{R}^{\kappa}\to\mathbb{R}^{\kappa}$ as $\Psi(x)=(p_{1}(x),...,p_{\kappa-1}(x),x_{\kappa})$. The Jacobian of $\Psi$ at 0 is then invertible. By the inverse function theorem, $\Psi^{-1}|_{U}=\Phi:U\to V$ is an analytic local diffeomorphism over neighborhoods $U$ and $V$ of 0. Let now $u=(u_{1},...,u_{\kappa-1})$ and define on $U$ the map $H(u,w)=p(\Phi(u,w))=(u,h(u,w))$, where $h$ is analytic by construction. The analytic diffeomorphism $\Phi$ now induces an isomorphism of $\mathbb{R}$-algebras $\mathbb{R}[[X_{1},....,X_{\kappa}]]/\langle p_{1},...,p_{\kappa}\rangle\cong\mathbb{R}[[u,w]]/\langle u,h(u,w)\rangle$ (one defines the pullback isomorphism $\Phi^{\#}:\mathbb{R}[[X_{1},...,X_{\kappa}]]\to\mathbb{R}[[u,w]]$ which takes $X_{i}\mapsto\Phi_{i}(u,w)$ and for each $F\in\mathbb{R}[[X_{1},...,X_{\kappa}]]$, $\Phi^{\#}(F)=F(\Phi(u,w))$. Computing the pullback on a generator $p_{i}$ gives $\Phi^{\#}(p_{i})=p_{i}(\Phi(u,w))$, which implies $\Phi^{\#}(I)=\langle u,h(u,w)\rangle$. Passing then the isomorphism to quotients gives the result. Note that the isomorphism constructed is also $\mathbb{R}$-linear, so it is a $\mathbb{R}$-algebra isomorphism).
Each class of $\mathbb{R}[[u,w]]/\langle u,h(u,w)\rangle$ can be represented by a power series in $w$ only, since each variable in $u$ is zeroed by the ideal $\langle u,h(u,w)\rangle$. This implies that we have an $\mathbb{R}$-algebra isomorphism $\mathbb{R}[[u,w]]/\langle h(0,w)\rangle\cong\mathbb{R}[[u,w]]/\langle u,h(u,w)\rangle$. Therefore we have obtained that
$\mathbb{R}[[X_{1},....,X_{\kappa}]]/\langle p_{1},...,p_{\kappa}\rangle\cong\mathbb{R}[[w]]/\langle g\rangle,$
where $g=h(0,\cdot)$ as $\mathbb{R}$-algebras. This concludes the proof of the“only if” part of the Lemma 3.5.
Now we prove the sufficiency of the Lemma 3.5. Suppose then $\mathbb{R}[[X_{1},....,X_{\kappa}]]/\langle p_{1},...,p_{\kappa}\rangle\cong\mathbb{R}[[w]]/\langle g(w)\rangle$. The algebra $Q=\mathbb{R}[[w]]/\langle g(w)\rangle$ has a unique maximal ideal given by $\mathfrak{M}=\langle w\rangle/\langle g(w)\rangle$. The set $\mathfrak{M}/\mathfrak{M}^{2}$ can be viewed as real vector space obtained as the quotient of the subspace $\mathfrak{M}$ by the subspace $\mathfrak{M}^{2}\subset\mathfrak{M}$. Its dimension is 0 or $1$.
The ring $\mathbb{R}[[X_{1},..,X_{\kappa}]]$ has a unique maximal ideal $\mathfrak{m}=\langle X_{1},...,X_{\kappa}\rangle$ and $A=\mathbb{R}[[X_{1},..,X_{\kappa}]]/I$ has a unique maximal ideal denoted $\mathfrak{m}_{A}=\mathfrak{m}/I$, with $I=\langle p_{1},...,p_{\kappa}\rangle$. Similarly to the above $\mathfrak{m}/\mathfrak{m}^{2}$ is a real vector space and the isomorphism between $A$ and $Q$ implies that $\mathfrak{m}_{A}/\mathfrak{m}^{2}_{A}\cong\mathfrak{M}/\mathfrak{M}^{2}$ as real vector spaces.
We can write $\mathfrak{m}^{2}_{A}=\frac{\mathfrak{m}^{2}+I}{I}$. Therefore,
$\mathfrak{m}_{A}/\mathfrak{m}^{2}_{A}\cong\frac{\mathfrak{m}/I}{(\mathfrak{m}^{2}+I)/I}\cong\frac{\mathfrak{m}}{\mathfrak{m}^{2}+I},$
as real vector spaces. Let $q:\mathfrak{m}\to\frac{\mathfrak{m}}{\mathfrak{m}^{2}}$ be a linear quotient map associating a power series to its linear part. Consider $\ell_{j}=(\frac{\partial p_{j}}{\partial x_{1}}(0),...,\frac{\partial p_{j}}{\partial x_{\kappa}}(0))$. Then $q(I)=\text{span}_{\mathbb{R}}\{\ell_{1},...,\ell_{\kappa}\}$. Therefore, we have the following isomorphisms of real vector spaces
$\frac{\mathfrak{m}}{\mathfrak{m}^{2}+I}\cong\frac{\frac{\mathfrak{m}}{\mathfrak{m}^{2}}}{q(I)}\cong\frac{\mathbb{R}^{\kappa}}{\text{span}_{\mathbb{R}}\{\ell_{1},...,\ell_{\kappa}\}}.$
The dimension of $\text{span}_{\mathbb{R}}\{\ell_{1},...,\ell_{\kappa}\}$ equals the rank of $Jp(0)$. Therefore, the dimension of $\frac{\mathbb{R}^{\kappa}}{\text{span}_{\mathbb{R}}\{\ell_{1},...,\ell_{\kappa}\}}$ is
$\kappa-\text{dim}_{\mathbb{R}}Jp(0)=\text{dim}_{\mathbb{R}}(\frac{\mathfrak{m}_{A}}{\mathfrak{m}^{2}_{A}}).$
Therefore, $\text{dim}_{\mathbb{R}}Jp(0)$ is either $\kappa-1$ or $\kappa$, showing that $(G,\sigma^{*})$ is monogenic. This concludes the proof of Lemma 3.5. ∎
We now prove $(b)$. Since it is known that the absence of payoff-robustness implies 0 index, we focus on showing the opposite direction. This will require the following auxiliary lemma:
Lemma 3.6.
Suppose that the pair $(G,\sigma^{*})$ is monogenic and let $g$ be the power series obtained in Lemma 3.5. If $\sigma^{*}$ has index zero then the order of $g$ is even. Conversely, if the order of $g$ is even, then $p$ is not locally surjective (and therefore $\sigma^{*}$ has index zero).
Proof of Lemma 3.6.
Suppose first that $\sigma^{*}$ has index zero. Proposition 2.1 then implies that $\text{dim}_{\mathbb{R}}\mathbb{R}[[w]]/\langle g(w)\rangle$ is even. Fixing the local order in $\mathbb{R}[[w]]$, a basis for $\mathbb{R}[[w]]/\langle g(w)\rangle$ comprises those monomials $w^{\alpha}$ which do not belong to the leading term ideal $\langle LT[\langle g(w)\rangle]\rangle$. Since this leading term ideal is generated by the term corresponding to ord($g$), say, $cw^{k}$, for some $c\neq 0$, it follows that a basis for $\mathbb{R}[[w]]/\langle g(w)\rangle$ is given by $\{1,....,w^{k-1}\}$, which then implies that $k$ is an even positive integer.
For the opposite direction, suppose that $g$’s order is even. Let $cw^{k}=\text{ord}(g)=\text{ord}(h(0,w))$ and suppose without loss of generality that $c>0$ (the proof for $c<0$ is entirely analogous). For a sequence $\delta_{k}\uparrow 0$, let $q_{k}=(0,\delta_{k})\in\mathbb{R}^{\kappa-1}\times\mathbb{R}$. Then $p(\Phi(u,w))=(0,\delta_{k})$ implies that $h(0,w)=\delta_{k}<0$. Take $U\subset\mathbb{R}^{\kappa-1}\times\mathbb{R}$ a neighborhood of 0 such that $h(0,w)>0$, for all $(0,w)\in U$. Such neighborhood exists because the order of $h(0,w)$ is even. Hence, there exists $k_{0}$ such that for each $k\geq k_{0}$, $(p\circ\Phi)^{-1}(q_{k})\cap U=\emptyset$. Therefore, $p\circ\Phi$ is not locally surjective at 0 and since $\Phi$ is a local diffeomorphism, $p$ is not locally surjective at 0, as well. This, together with 0 being a locally isolated root of $p$, implies that $\text{deg}_{0}(p)=0$. Proposition 3.1 implies then that the index of $\sigma^{*}$ equals 0. This concludes the proof of the lemma. ∎
We now conclude the proof of part $(b)$. Let $\hat{J}$ be the ideal defined by the following set of equations in $Z_{n}=(Z_{s^{0}_{n}},...,Z_{s^{k_{n}}_{n}})$ and $Y_{n}=(Y_{s^{0}_{n}},....,Y_{s^{k_{n}}_{n}})$ (below, $\mathbbm{1}$ denotes the vector with $1$ in all coordinates),
$Z_{n}-Y_{n}-G^{n}(Y)=0,\forall n\in\mathcal{N}$
$(Z_{n}-Y_{n})\cdot(e_{s_{n}}-Y_{n})=0,\forall s_{n}\in S_{n},n\in\mathcal{N}$
$Y_{n}\cdot\mathbbm{1}-1=0,n\in\mathcal{N}$
Let $\hat{z}$ be the solution to $(\text{proj}\circ\theta^{-1}_{G})(z)=0$, for which $r_{n}(\hat{z}_{n})=\sigma^{*}_{n}$ for all $n\in\mathcal{N}$. Let us change the variable $Y_{n}$ to $Y_{n}+\sigma^{*}_{n}$ and $Z_{n}$ to $Z_{n}+\hat{z}_{n}$ so that the isolated completely mixed solution $Y_{n}=r_{n}(\hat{z}_{n})$ and $Z_{n}=\hat{z}_{n}$ is translated to zero. Under this change of variables, the ideal generated by the set of equations is denoted by $J$. Now, since $(\text{proj}\circ\theta^{-1}_{G})(z)$ is a polynomial in a neighborhood of $\hat{z}$ and considering the change of variables from $Z_{n}$ to $Z_{n}+\hat{z}_{n}$, denote by $T(Z)$ this polynomial after the change of variables. The proof of part $(b)$ will follow from verifying that the sequence of $\mathbb{R}$-algebra isomorphisms holds:
$\frac{\mathbb{R}[[Y]]}{\langle p_{1},...,p_{\kappa}\rangle}\cong\frac{\mathbb{R}[[Z,Y]]}{J}\cong\frac{\mathbb{R}[[Z]]}{\langle T(Z)\rangle}.$
Assume first that the sequence of isomorphisms above holds in order to conclude the proof. Because $(G,\sigma^{*}$) is monogenic and $\sigma^{*}$ has index 0, the Jacobian $Jp(0)$ must lose full rank by exactly $1$. The chain of isomorphisms above then implies that the Jacobian of $T$ at 0 loses rank by exactly $1$ as well. Applying the same reasoning we applied to $p$ in the proof of Lemma 3.6, we obtain that $T$ is not locally surjective at 0. Therefore, $\text{proj}\circ\theta^{-1}_{G}$ is not locally surjective at $\hat{z}$, which is equivalent to $\hat{z}$ not being payoff-robust, which implies our desired result.
We now construct the claimed sequence of isomorphisms: $\varphi:\mathbb{R}[[Z]]\to\mathbb{R}[[Z,Y]]/J$ taking $Z\mapsto[Z]$. Note that the relations in $J$ determine that $Y_{n}+\sigma^{*}_{n}$ is the retraction of $Z_{n}+\hat{z}_{n}$ to the affine space generated by player $n$’s strategy, when $(Z_{n},Y_{n})$ takes values close to zero. As this retraction can be expressed as a polynomial in the variables of $Z_{n}$, it implies that each $Y_{n}$ can be substituted by a polynomial $q_{n}(Z_{n})$ in the variables $Z_{n}$ in the algebra $\mathbb{R}[[Z,Y]]/J$. Therefore, each equivalence class has a representative in $\mathbb{R}[[Z]]$. Therefore, $\varphi$ is surjective. Now, the kernel of $\varphi$ is simple to compute: note that $(\text{proj}\circ\theta^{-1}_{G})_{n}(z)=z_{n}-r_{n}(z_{n})-G^{n}(r_{-n}(z_{-n}))$. Therefore, substituting $Y_{n}=q_{n}(Z_{n})$ in $J$, the ideal $J$ reduces to $\langle T(Z)\rangle$. Therefore, the kernel of $\varphi$ is $\langle T(Z)\rangle$, which proves the second isomorphism, by passing $\varphi$ to the quotient. Note that the isomorphism is $\mathbb{R}$-linear.
For the first isomorphism, note that the relations in $J$ allow us to eliminate $Z$ in $\mathbb{R}[[Z,Y]]/J$ by isolating $Z_{n}$ in the first set of equations. Substituting the expression for $Z_{n}$ in the second set of equations gives us precisely the polynomials $p_{1},....,p_{\kappa}$ (with the third set of equations being used to eliminate one strategy variable per player). The implied isomorphism is also $\mathbb{R}$-linear. This concludes the proof of the sequence of isomorphisms and the proof of the theorem.∎
Remark 3.7.
As can be seen from the proof of item $(a)$ of Theorem 3.4, this makes no use of the particular fact that the polynomials originate from equilibrium conditions in a finite game. As long as 0 is isolated as a root of a polynomial map $p$, it holds. Obviously, this is not the case with item $(b)$, which makes use of the particular polynomials describing completely mixed equilibria of the finite game.
Corollary 3.8.
Suppose $(G,\sigma^{*})$ is monogenic. The index of $\sigma^{*}$ is zero if and only if the dimension of the $\mathbb{R}$-algebra $\mathbb{R}[[X_{1},....,X_{\kappa}]]/\langle p_{1},...,p_{\kappa}\rangle$ is even.
Proof.
Follows from the formula in Proposition 2.1 and item (a) of Theorem 3.4. ∎
As the corollary suggests, losing full rank in the monogenic case is not sufficient to guarantee that the index is zero, i.e., it is possible that a completely mixed and isolated equilibrium of a game has an associated Jacobian without full rank and yet has index different than 0, as Proposition 3.9 shows. This is a consequence of the proof of Theorem $7$ in Datta (2003).
Moreover, when there are two players, it is known that an isolated completely mixed equilibrium implies that the Jacobian has full rank: in this case, there are no completely mixed isolated equilibria with zero index, and consequently any isolated completely mixed equilibrium is payoff-robust. The same is not true when there are more players, and one can construct (monogenic) examples where an isolated completely mixed equilibrium has a singular Jacobian with index 0 (which by (b) in Theorem 3.4 implies it is not payoff-robust).
Proposition 3.9.
There exists a finite game with three players where one player has $2$ strategies and the other two have $3$ strategies each, with a completely mixed equilibrium $\sigma^{*}$ such that $(G,\sigma^{*})$ is monogenic with rank of $Jp(0)$ equal to $\kappa-1$ and the index is nonzero.
In order to make the proof of Proposition 3.9 more self-contained we make a few observations on the method used by Datta (ibid.). The proposition relies on a simple way of writing a one-variable polynomial system as a logically equivalent system of polynomial equations that ultimately describe the completely mixed equilibrium of a game. If $q(x)=\alpha_{d}x^{d}+\alpha_{d-1}x^{d-1}+...+\alpha_{0}=0$ is a one variable polynomial equation, then it can be rewritten iteratively as $q(x)=(...(\alpha_{d}x+\alpha_{d-1})x+\alpha_{d-2})x+...+\alpha_{1})x+\alpha_{0}=0$. We can then “disaggregate” this polynomial equation into a system of multiaffine equations in two variables $c_{i}^{\prime}s$ and $x$: for example, one could write:
(2) $0$ $=\alpha_{0}+x(s_{1}c_{1}+\delta_{1})$
$s_{1}c_{1}+\delta_{1}$ $=\alpha_{1}+x(s_{2}c_{2}+\delta_{2})$
$\vdots$
$s_{d-1}c_{d-1}+\delta_{d-1}$ $=\alpha_{d-1}+x\alpha_{d}$
Note that each such equation is multiaffine in the variables $c_{i}$’s and $x$, and $s_{i}^{\prime}s$ and $\delta_{i}$’s are constants chosen so that the solutions in $c_{i}$ end up being probabilities (supposing that $q$ has a solution $x\in(0,1)$). Lemma 4 in Datta (ibid.) shows that systems such as the one above correspond to some of the polynomial equations describing the completely mixed equilibria of a game, namely, the payoff indifference equations of a third player (one that does not have the strategies $c_{i}$ or $x$).
Proof of Proposition 3.9.
The proof is a corollary of the proof of Theorem 7 in Datta (2003), so we simply sketch the reasoning for completeness, using the notation of that theorem. Consider a real polynomial in one-variable $q(a)$ such that $q(0)=0$, the degree of $q$ is $4$ and the order of $q$ is $3$. Fix some $a^{*}\in(0,1)$ and consider then the translated polynomial $p(a)=q(a-a^{*})$. The polynomial $p$ has an isolated root at $a^{*}$. By Theorem 7 in Datta (2003), there exists a three-player (Alice, Bob and Critter) game $G$ where player Alice has 2 pure strategies, and Bob and Critter have $3$ strategies each. Consider the polynomial system $Q(a,b_{1},b_{2},c_{1},c_{2})=0$ defining the completely mixed equilibria of the game as outlined in p. 430, ibid. (we follow the notation used in Datta’s proof, where $a$ denotes one of the mixed strategy coordinates of Alice, $b_{1}$, $b_{2}$ denote two of the mixed strategy coordinates of Bob, and $c_{1},c_{2}$ two mixed strategy coordinates of Critter). By construction of the system, there exists a uniquely determined completely mixed equilibrium $\sigma^{*}=(a^{*},b^{*}_{1},b^{*}_{2},c^{*}_{1},c^{*}_{2})$ (here we omitted a strategy variable for each player, since it is uniquely determined by probabilities) by the root $a^{*}$. Now, change variables in the system $Q=0$, so that the completely mixed equilibrium $\sigma^{*}$ is translated to 0 in a new system $F(A,B_{1},B_{2},C_{1},C_{2})=0$ with an isolated root at 0. The system can now be reduced, through a series of substitutions and eliminations of variables (see p. 427, ibid. for an explanation), to $q(a)=0$, thus implying that $\mathbb{R}[[A,B_{1},B_{2},C_{1},C_{2}]]/\langle F\rangle\cong\mathbb{R}[[a]]/\langle q(a)\rangle$, as $\mathbb{R}$-algebras. Therefore $(G,\sigma^{*})$ is monogenic (by Lemma 3.5). Moreover, the Jacobian $JF(0)$ loses rank by $1$, since the maximal ideal of $\mathbb{R}[[a]]/\langle q(a)\rangle$ is non-zero (the elements of the ideal $\langle q(a)\rangle$ all have order at least $3$, so the maximal ideal $\mathfrak{M}=\langle a\rangle/\langle q(a)\rangle$ is non-zero. This in turn implies that $\mathfrak{M}/\mathfrak{M}^{2}$ has dimension $1$, which determines that $JF(0)$ does not have full rank). Finally, recall that the dimension of the $\mathbb{R}$-algebra $\mathbb{R}[[x]]/\langle q(x)\rangle$ is odd, since ord(q)=3(q)=3, which implies that the index is nonzero (by Lemma 3.6).∎
Remark 3.10.
Note that, since the proof of Theorem $7$ in Datta (2003) is constructive, one could obtain the game of Proposition 3.9 explicitly.
Example 3.11.
As an illustration of the results of Theorem 3.4, we revisit the example in subsection 1.1 This is a three-player game with two strategies per player (named $a$ and $b$), with the standard matrix representation below:
$a$ $b$
$a$ $(1,1,1)$ (5,0,3)(-5,0,3)
$b$ (0,3,5)(0,3,-5) $(0,0,1)$
Player 3 chooses $a$
$a$ $b$
$a$ (3,5,0)(3,-5,0) $(1,0,0)$
$b$ $(0,1,0)$ $(0,0,0)$
Player 3 chooses $b$
This game has a completely mixed equilibrium $\sigma^{*}$ where each player randomizes with equal probabilities between $a$ and $b$. From this we can compute the polynomial map $p$ corresponding to this completely mixed equilibrium: letting $x$ denote the probability that player 1 puts in $a$, $y$ the probability that player 2 puts in $a$, and $z$ the probability that player 3 puts in $a$, we have the following system of equations:
(3) $p_{1}$ $=x-y+xy=0$
$p_{2}$ $=y-z+yz=0$
$p_{3}$ $=z-x+zx=0$
The equilibrium $\sigma^{*}$ is isolated and a quick verification of the Jacobian of the system gives that $(G,\sigma^{*})$ is monogenic. In order to verify that $\sigma^{*}$ has index 0, we use the fact that the algebra given by $\mathbb{R}[[x,y,z]]/\langle p_{1},p_{2},p_{3}\rangle$ has the same dimension as $Q=\mathbb{R}[[x,y,z]]/\langle LT[\langle p_{1},p_{2},p_{3}\rangle]\rangle$, and the latter is immediate to compute: fix the degree anticompatible lex order given by $x>y>z$: then, immediately from $p_{1}$ and $p_{2}$, $x\in LT[\langle p_{1},p_{2},p_{3}\rangle]$, $y\in LT[\langle p_{1},p_{2},p_{3}\rangle]$. Therefore, a basis for $Q$ is given by $\{1,z,...,z^{k}\}$, for some non-negative integer $k$. We only need to verify now that $k$ is odd in order to obtain that $\sigma^{*}$ has index 0 (if $k$ is even, then it has index +1+1 or 1-1). We can verify this by simple substitution: from $p_{2}$, y(1+z)z=0y(1+z)-z=0, yielding $y=\frac{z}{1+z}$. From $p_{3}$, x(z1)+z=0x(z-1)+z=0, yielding $x=\frac{z}{1-z}$. Now substituting in $p_{1}$, it yields $\frac{3z^{2}}{(1-z)(1+z)}$, and since $\frac{3}{(1-z)(1+z)}$ is a unit in $\mathbb{R}[[x,y,z]]$, it follows that $\{1,z\}$ is a basis for $Q$, since $z^{2}\in LT[\langle p_{1},p_{2},p_{3}\rangle]$. As observed in Corollary 3.8, because the dimension is even, it follows that the index of $\sigma^{*}$ is 0. From part $(b)$ of Theorem 3.4, it follows that $\sigma^{*}$ is not payoff-robust. In p.151 of McLennan (2018), a specific payoff perturbation is constructed to show that $\sigma^{*}$ is not robust. Theorem 3.4 allows us to bypass this construction and simply check the dimension of $Q$.
Generalizing from the analysis in the last example, we have the more general statement about three-player games with $2$ strategies per player.
Proposition 3.12.
Let $G$ be a finite, three-player game with $2$ strategies per player. Let $\sigma^{*}$ be a completely mixed equilibrium of $G$ that is isolated. Then $(G,\sigma^{*})$ is monogenic. Therefore, the indices of isolated equilibria in such games can only be 0, 1-1 and +1+1.
Proof.
Let $p=p_{1}\times p_{2}\times p_{3}$ be the polynomial map associated to $(G,\sigma^{*})$ in variables $x$ (player $1$’s strategy), $y$ (player $2$’s strategy) and $z$ (player $3$’s strategy). We view $p_{1}$ as the polynomial of player $1$, that is, the polynomial featuring variables corresponding to strategies of players $2$ and $3$ only. Analogously, we view $p_{2}$ as the polynomial of player $2$ and $p_{3}$ as the polynomial of player $3$. Because 0 is an isolated root of $p$ by assumption, the linear parts of $p_{1},p_{2}$, and $p_{3}$ must be such that they feature linear monomials in $x,y,$ and $z$: if not, by the multilinearity of each $p_{i}$, the absence of such a monomial in the linear part of $p_{i}$ for each $i$ implies that 0 is not isolated. Now, in each $p_{i}$, there is at least one variable absent, namely, the variable corresponding to the strategy of player $i$. Let $\ell(p_{1})$, $\ell(p_{2})$ and $\ell(p_{3})$ be the linear parts of polynomials $p_{1},p_{2}$, and $p_{3}$, respectively. Therefore, we have the following possibilities: (1) each $\ell(p_{i})$ features at least one variable; (2) exactly one $\ell(p_{i})=0$, with the others $\ell(p_{j}),j\neq i$ featuring all the remaining variables.
In the first case, suppose without loss of generality that $\ell(p_{2})$ has a linear term in $x$. Then $y$ or $z$ must feature in $\ell(p_{1})$. Without loss of generality, suppose that it is $y$. Fix the degree anticompatible lex order >>, with $x>y>z$. Construct the $3\times 3$ matrix $M$ as follows: entry $M_{ij}$ is the coefficient of the variable of lex-order $j$ in equation $i$. As the linear term in $x$ cannot feature in $\ell(p_{1})$, $M_{11}=0$. Similarly, $M_{22}=M_{33}=0$. Since $y$ features in $\ell(p_{1})$, then $M_{12}\neq 0$. We then have that the rank of $M$ is at least $2$ (for example, by Gaussian elimination). This implies that $(G,\sigma^{*})$ is monogenic.
In the second case, suppose without loss of generality that $\ell(p_{1})=0$, and $\ell(p_{2})$ features two variables, with $\ell(p_{3})$ featuring exactly one. Again without loss of generality, suppose that $x$ features in $\ell(p_{2})$ and $z$ features in $\ell(p_{3})$. Fixing the degree anticompatible lex order >>, with $x>z>y$, and constructing $M$ as above, implies, by the exact same reasoning, that $M$ has rank at least $2$, which implies that $(G,\sigma^{*})$ is monogenic.∎
A natural question to ask now is what happens with the indices of isolated and completely mixed equilibria outside the class of monogenic games. The answer to this question is given in Theorem 3.13. This theorem can be viewed as a strengthening of the result in Proposition 3.1 in Govindan et al. (2004), which proved that for any integer $n$, one can obtain a (bi-matrix) game with a nondegenerate component of equilibria of index $n$. Theorem 3.13 can also be viewed as the game-theoretic analog of the known fact of fixed point theory: for any integer $n$, it is possible to show that there exists a map from $\mathbb{R}^{2}$ to $\mathbb{R}^{2}$ with an isolated fixed point at 0 with index $n$.
Theorem 3.13.
Given any integer $n\in\mathbb{Z}$, there exists a three-player game $G$ with an isolated completely mixed equilibrium with index $n$.
Proof.
We first construct polynomials with an isolated root at 0 with any fixed degree. Suppose $m>0$. Consider the complex function $z=(x+iy)\mapsto(x-iy)^{m}$. Let $F:\mathbb{R}^{2}\to\mathbb{R}^{2}$, be the realization of the complex function as a real polynomial map. The polynomial map $F$ has a unique root at 0. By construction the winding number of $F$ around 0 is $m$, thus implying that $\text{deg}_{0}(F)=-m$. Alternatively, consider $z\mapsto z^{m}$ and the realization $F$ of this polynomial to obtain the unique root at 0 with topological degree $m$. Now suppose $m=0$. Consider the real polynomial map $F(x,y)=(x^{2},y)$. This polynomial has an isolated root at 0. We claim the topological degree of 0 w.r.t. $F$ is 0: first we note that the dimension of the $\mathbb{R}$-algebra $\mathbb{R}[[X,Y]]/\langle X^{2},Y\rangle$ is $2$, since $\{1,X\}$ are the only surviving monomials under the relations imposed by the ideal (and $\{1,X\}$ is a basis for the space). Therefore, the formula for the computation of the absolute value of the topological degree in Proposition 2.1 applies. Now note that the maximal square-zero ideal of $\mathbb{R}[[X,Y]]/\langle X^{2},Y\rangle$ is generated by $X$. It is, therefore, one-dimensional. The formula then implies that 0 has topological degree 0.
In order to apply Theorem 5 in Datta (2003) and obtain a finite game associated to the zero-set of each of the polynomials above, we change variables so that the unique root 0 is translated to the point where all coordinates are equal to $1/3$, satisfying the requirement of that theorem. Of course, the topological degree of 0 in either case is invariant under this translation. Applying Theorem 5, we obtain the required game. Now, one needs to verify that the topological degrees of the equilibria are maintained under the new system describing the completely mixed equilibria. For that, consider $F$ as one the polynomial maps constructed above. We checked that in the case of $F(x,y)=(x^{2},y)$, the algebra $\mathbb{R}[[X,Y]]/\langle X^{2},Y\rangle$ is finite dimensional, and a similar argument yields the same finite-dimensionality for $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$ in the case $F$ is specified as the other polynomials. Applying Theorem 5 in Datta, this implies that the absolute value of the topological degree of the equilibria is maintained, but one needs to show that the signs are maintained as well, so we need a stronger result. We use Theorem 1.2 in Eisenbud et al. (1977), which provides a formula for the topological degree based on the signature of the bilinear form defined over the $\mathbb{R}$-algebra. Proposition 2 in Eisenbud (1978) yields that the minimal non-zero ideal in $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$ is one-dimensional and is generated by the class of the Jacobian determinant of this polynomial map. Since $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$ is isomorphic to $\mathbb{R}[[X_{1},...,X_{\kappa}]]/\langle p_{1},...,p_{\kappa}\rangle$ (the $\mathbb{R}$-algebra associated to the system of completely mixed equilibria associated to $F$ and obtained by Theorem 5 in Datta), that minimal non-zero ideal maps to the minimal non-zero ideal of $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$ under the isomorphism. Call this isomorphism $\varphi$. Since the minimal non-zero ideals of $\mathbb{R}[[X_{1},...,X_{\kappa}]]/\langle p_{1},...,p_{\kappa}\rangle$ and $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$ are one-dimensional, $\varphi$ maps one to the other, up to a constant $c$. In the particular case of the substitutions implied by Datta’s argument in the proof of Theorem 5, we can take $c=1$. This implies that if $\phi:\mathbb{R}[[X_{1},...,X_{\kappa}]]/\langle p_{1},...,p_{\kappa}\rangle\to\mathbb{R}$ is a linear functional that takes value $1$ in the class of the Jacobian determinant, then the signature of the bilinear form $a,b]:=\phi(ab),a,b\in\mathbb{R}[[X_{1},...,X_{\kappa}]]/\langle p_{1},...,p_{\kappa}\rangle$ is equal to the topological degree of 0 under $p$. The isomorphism $\varphi$ allows us to define the linear functional $\psi=\phi\circ\varphi^{-1}$ on $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$ that takes value $1$ on the class of the Jacobian determinant of $\mathbb{R}[[X,Y]]/\langle F_{1},F_{2}\rangle$. Note that the signaure of $\cdot$$\cdot$] equals the signature of $(\cdot,\cdot)$, by construction. By Theorem 1.2 in Eisenbud and Levine (ibid.), the signature of $(\cdot,\cdot)$ is equal to the topological degree of 0 under $F$. The proof is then concluded by observing that the topological degree and index of the equilibrium point agree, by Proposition 3.1. ∎
Remark 3.14.
In a brief comment in section 4 of Datta (2003), the author observes that, by viewing the map $z\mapsto z^{n}$ as a map in $\mathbb{R}^{2}$ one is able to generate a two-variable real polynomial map $F$ with 0 as an isolated root, with topological degree $n\in\mathbb{N}$, which implies from the results of the paper that any natural number could be the topological degree of a completely mixed equilibrium. The author here does not seem to refer to the canonical notion of topological degree of equilibrium derived from the Structure Theorem (Kohlberg and Mertens (1986)), but simply to the topological degree determined by the polynomial $F$ to 0.
4. Extensions
There are straightforward extensions of the index formula presented in Proposition 2.1. First, the index of equilibria is invariant under the elimination of duplicate strategies (i.e., mixed strategies of a player that are added to the game as pure strategies of that player). This implies that if there is a non-degenerate component of equilibria of a game that reduces to a singleton completely mixed equilibrium once duplicate strategies are eliminated, then this implies that the indices of the component and the equivalent singleton are the same. As an example, if we add to the three-player game in Example 3.11 the completely mixed equilibrium mixed strategy of player 1 as a pure strategy, then a nondegenerate component of equilibria arises, and this component has the same index as the singleton once the duplicate strategy is eliminated. Therefore, if the singleton is monogenic in the reduced game, then the results of Theorem 3.4 apply as well.
Second, the index is also invariant under the addition or elimination of strictly inferior replies to the equilibrium: therefore, if a singleton equilibrium lies on the boundary of the strategy set and after elimination of inferior replies, it becomes completely mixed in the reduced game, then the results of this paper apply to such boundary cases. As an illustration, consider the following outside-option game with three players: player 1 first decides between the outside-option (where the game ends and players get (1,2,7)(-1,2,7)) or coming In and playing the simultaneous move game defined in Example 3.11. Notice that the payoff to player 1 of the equilibrium where player 1 chooses In and then the completely mixed equilibrium ($\sigma^{*}$) is played is 0 (this equilibrium is on the boundary of the mixed strategy set). Any strategy of player 1 where $Out$ is chosen first is a strictly inferior reply to that equilibrium. Therefore, the index of the equilibrium In-$\sigma^{*}$ is 0 in the outside-option game.
A third (less straightforward) extension of the results in this paper concerns extensive form games with $N$ players, namely, equilibrium outcomes that are isolated and completely mixed (i.e., where the distribution over terminal nodes induced by equilibria is completely mixed and isolated in the space of equilibrium distributions). [^5] For a generic choice of terminal payoffs in a fixed game tree with perfect recall, one has that each one of the finitely many connected components in mixed strategies of the game yields a unique equilibrium outcome distribution. In order to compute the index of an isolated completely mixed equilibrium outcome, one first needs to compute the strategies inducing the outcomes in enabling form (see Govindan and Wilson (2002) or von Stengel (1996)): since the outcome is isolated and completely mixed, there is a unique enabling strategy profile $e^{*}$ inducing it, which is in the interior of the enabling strategy set of the players and is isolated in the space of equilibria in enabling form. In each enabling strategy space $E_{n}$ of player $n$, consider a simplex $\Delta_{n}$ satisfying the following: (1) $\Delta_{n}$ is contained in the interior of $E_{n}$; (2) $\Delta_{n}$ has the same dimension as $E_{n}$; (3) $\Delta_{n}$ contains $e^{*}_{n}$ in its interior; (4) $\Delta=\times_{n}\Delta_{n}$ contains no other equilibrium of the game besides $e^{*}$. Consider now the enabling form of this game (see section 4 of Pahl (2023)): if $V_{n}:\times_{n}E_{n}\to\mathbb{R}$ is player $n$’s multiaffine payoff function in enabling strategies, let $G_{n}(v_{1},...,v_{N}):=V_{n}(v_{1},...,v_{N})$, where $v_{n}$ is a vertex of $\Delta_{n}$. We can view each vertex $v_{n}$ as a pure strategy of player $n$ and therefore define a normal form game $G$ by considering the mixed extension over these pure strategies. This normal form game now has a completely mixed equilibrium $\sigma^{*}$ (which corresponds to $e^{*}$), and the index of $\sigma^{*}$ can be computed through the formula presented in Proposition 2.1. This index of the component of equilibria in mixed strategies that induces the equilibrium outcome distribution is identical to the normal form index of $\sigma^{*}$ in $G$ (see Proposition 3.10 in Pahl (2023)). This procedure outlines a method to compute the index of an isolated completely mixed equilibrium outcome of an extensive form game by reducing the initial game to a normal form game with an associated isolated completely mixed equilibrium that has the same index. The procedure described above in particular implies that in two-player extensive form games completely mixed isolated equilibrium outcomes have indices +1+1 or 1-1 only. For more players, if the completely mixed outcomes reduce to monogenic equilibria in enabling strategies, then 0, 1-1 or +1+1 are the only possible indices. The extension outlined above yields a “perturbation-free” algorithm to compute the index of isolated completely mixed outcomes.
Example 4.1.
We illustrate the method described above in the following 3-player extensive form game.
[…]
Let $x$ be the probability of $L$, $y$ be the probability of $U$ after $L$, $z$ the probability of $U$ after $R$, $w$ the probability of $A$ in the information set $I_{U}$ of player 3, and $v$ the probability of $A$ in the information set $I_{D}$ of player 3. This game has a completely mixed, isolated equilibrium given by:
$x^{\ast}=\frac{1}{2},\qquad y^{\ast}=\frac{1}{3},\qquad z^{\ast}=\frac{2}{3},\qquad w^{\ast}=\frac{1}{4},\qquad v^{\ast}=\frac{3}{4}.$
Because in this game no player can exclude an information set of herself by chosing any of her actions, the set of behavior strategies of each player is equal to her set of enabling strategies. Players 1’s enabling strategy set is equal to $E_{1}=[0,1]$, Player $2$’s enabling strategy set is equal to $E_{2}=[0,1]^{2}$ (with coordinates $(y,z)$) and Player $3$’s strategy set is equal to $E_{3}=[0,1]^{2}$ (with coordinates $(w,v)$). The multiaffine payoff functions of each player are given by:
$V_{1}(x,y,z,w,v)$ $=xy+(1-x)(1-z),$
$V_{2}(x,y,z,w,v)$ $=x\big(yw+(1-y)(1-v)\big)+(1-x)\big(zw+(1-z)\tfrac{v}{3}\big),$
$V_{3}(x,y,z,w,v)$ $=x\big(y(2w)+(1-y)v\big)+(1-x)\big(z(-w)+(1-z)(-2v)\big).$
We define now the auxiliary normal form game $\hat{G}$ to this enabling form in order to compute the index of that equilibrium. Fix $\varepsilon=0.05$. Define a 1-simplex for Player 1:
$\Delta_{1}=\mathrm{conv}\{\,x^{\ast}-\varepsilon,\ x^{\ast}+\varepsilon\,\}\subset(0,1).$
Define 2-simplices for Players 2 and 3:
$\Delta_{2}$ $=\mathrm{conv}\Big\{(y^{\ast}-\varepsilon,z^{\ast}-\varepsilon),\ (y^{\ast}+\varepsilon,z^{\ast}-\varepsilon),\ (y^{\ast},z^{\ast}+\varepsilon)\Big\}\subset(0,1)^{2},$
$\Delta_{3}$ $=\mathrm{conv}\Big\{(w^{\ast}-\varepsilon,v^{\ast}-\varepsilon),\ (w^{\ast}+\varepsilon,v^{\ast}-\varepsilon),\ (w^{\ast},v^{\ast}+\varepsilon)\Big\}\subset(0,1)^{2}.$
$e^{\ast}=(x^{*},y^{*},z^{*},w^{*},v^{*})$ lies in the interior of each $\Delta_{n}$, and $\Delta:=\Delta_{1}\times\Delta_{2}\times\Delta_{3}$ contains no other equilibrium besides $e^{\ast}$. Let the vertices of $\Delta_{2}$ be $v_{2}^{1},v_{2}^{2},v_{2}^{3}$ as listed above, and similarly for $\Delta_{3}$. Define a normal-form game $\hat{G}$ where Player $n$’s pure strategies are the vertices of $\Delta_{n}$, and payoffs are given as follows:
$\hat{G}_{n}(v_{1}^{i},v_{2}^{j},v_{3}^{k}):=V_{n}(v_{1}^{i},v_{2}^{j},v_{3}^{k}).$
Viewing vertices as pure strategies of the players in $\hat{G}$, this defines a normal form game with an isolated completely mixed equilibrium $\sigma^{*}$ corresponding to $e^{*}$ with the same index. The equilibrium $\sigma^{*}$ is given by writing $e^{*}$ in the barycentric coordinates of the simplices $\Delta_{n}$:
$x^{\ast}=\tfrac{1}{2}v_{1}^{1}+\tfrac{1}{2}v_{1}^{2},\qquad(y^{\ast},z^{\ast})=\tfrac{1}{4}v_{2}^{1}+\tfrac{1}{4}v_{2}^{2}+\tfrac{1}{2}v_{2}^{3},\qquad(w^{\ast},v^{\ast})=\tfrac{1}{4}v_{3}^{1}+\tfrac{1}{4}v_{3}^{2}+\tfrac{1}{2}v_{3}^{3}.$
We know employ the techniques developed in this paper to compute the index of $\sigma^{*}$. We first compute the polynomial system given by $p$ defined in the beginning of section 3. There are five equations in five unknowns $(P,Q_{1},Q_{2},R,S)$:
(4) $p_{1}$ $=3Q_{1}+Q_{2}=0$
(5) $p_{2}$ $=11R-7S+28\varepsilon PR+52\varepsilon PS=0,$
(6) $p_{3}$ $=-7R-13S-44\varepsilon PR+28\varepsilon PS=0$
(7) $p_{4}$ $=-24\varepsilon PQ_{1}-24\varepsilon PQ_{2}-16P+6Q_{1}+6Q_{2}=0$
(8) $p_{5}$ $=-168\varepsilon PQ_{1}-72\varepsilon PQ_{2}-16P+18Q_{1}+42Q_{2}=0$
We compute the topological degree of the root $(P,Q_{1},Q_{2},R,S)=(0,0,0,0,0)$ using the formula of Proposition 2.1, which then gives us the index of that the completely mixed equilibrium $\sigma^{*}$, which in turn gives us the index of the connected component in mixed strategies inducing the outcome.
First use equation (4) to eliminate $Q_{2}$ and substitute the result in (7) and (8), thus obtaining a system in $4$ variables and $4$ equations. Equations (7) and (8) now become:
(9) $48\varepsilon PQ_{1}-16P-12Q_{1}=0,$
(10) $48\varepsilon PQ_{1}-16P-108Q_{1}=0.$
Lets fix the degree anticompatile lex order given by $Q_{1}<P<R<S$. Note now that a linear combination of the equations (9) and (10) (which is of course in the ideal $I$ generated by the polynomial map $p$) gives an equation in one variable only ($Q_{1}$). Therefore $Q_{1}\in LT[I]$. In addition, from equation (9), we have that $P$ also belongs to $LT[I]$. From equation (5), $S$ also belongs to $LT[I]$. Therefore, $\mathbb{R}[[Q_{1},P,R,S]]/LT[I]$ has a basis of monomials in $R$ only. But using the combination of (9) and (10), one gets that $Q_{1}=0$. Substituting this equation in (9), we get $P=0$ as well. Substituting then the latter in (5) and (6), we get two linear equations in $R$ and $S$. Eliminating $S$ from these equations, we get a linear equation in $R$, which shows that $R\in LT[I]$. Therefore, $\mathbb{R}[[Q_{1},P,R,S]]/LT[I]$ has dimension $1$ with $I_{max}=0$. Therefore, the absolute value of the degree is $1$.
5. Final Remarks
There are important cases that do not necessarily fit the framework presented in this paper: generic extensive-form games frequently involve non-degenerate components of equilibria that are located on the boundary of the strategy set (in mixed strategies). For example, in the case of two-player outside-option games, where the outside-option is an equilibrium. In such cases, even under the transformations we discussed in the previous paragraph, one cannot in general reduce the equilibrium component to a completely mixed equilibrium after the elimination of duplicates or inferior replies, or recasting the game in enabling form. These examples indicate that more work needs to be done in order to extend these algebraic methods to non-degenerate components of equilibria which are possibly not completely mixed.
Finally, there is an important question about the relation between non-zero index and payoff-robustness that this paper was not able to answer: are there completely mixed equilibria that have index 0, but are payoff-robust? At first, it seemed to us that this question could be handled by the methods presented in Datta (2003) in combination with the algebraic methods developed here: Datta’s method allows us to reduce questions about equilibria in games to questions about real algebraic varieties, but this reduction does not seem to preserve payoff-robustness of equilibria: a zero of a polynomial map might be robust to perturbations of the coefficients of the polynomial map, but the corresponding equilibrium to that zero - after applying Datta’s methods - might not be payoff-robust. The issue arises because the polynomial systems constructed in Datta that describe the completely mixed equilibria of a game and reduce to the original polynomial map after substitutions are highly non-generic, a fact that yields a payoff space that is possibly higher dimensional than the space of coefficients of the original polynomial map.